OpenStudy (brokenstanzas):

If the inverses of two functions are both functions, will the inverse of the sum or difference of the original functions also be a function? I need to provide examples also

satellite73 (satellite73):

the question is this: if \(f\) and \(g\) are one to one functions, are \(f+g\) and \(f-g\) also one to one ? if so, they will have an inverse that is a function

satellite73 (satellite73):

i am not sure what was written above, but here is an example to think about suppose \[f(x)=x+1\] amd \[g(x)=-x+1\] both are one to one, they are lines but what is \[(f+g)(x)\]?

OpenStudy (mathmate):

Or think about f(x)=x^3 g(x)=-x (f+g)(x)=x^3-x and decide what you can conclude.